The first step in HNO3 production is the catalyzed oxidation


of NH3. Without a catalyst, a different reaction predominates:  


4NH3(g) + 3O2(g) ⥫⥬ 2N2(g) + 6H2O(g)


When 0.0150 mol of NH3(g) and 0.0150 mol of O2(g) are placed


in a 1.00-L container at a certain temperature, the N2 concentra-


tion at equilibrium is 1.96×10−3 M. Calculate Kc

1
Expert's answer
2021-11-09T10:28:28-0500

Initial concentration = moles/volume = moles/1.00 L = initial moles


4 NH3(g) + 3 O2(g) <=> 2 N2(g) + 6 H2O(g)


I 0.0150 0.0150 0 0


C -4a -3a +2a +6a


E 0.0150-4a 0.0150-3a 2a 6a

[N2] = 2a = 1.96 x 10-3 = 0.00196 M

a = 0.00098

[NH3] = 0.0150 - 4a = 0.01108 M

[O2] = 0.0150 - 3a = 0.01206 M

[H2O] = 6a = 0.00588 M

Kc = [N2]2[H2O]6/[NH3]4[O2]3

= 0.001962 x 0.005886/(0.011084 x 0.012063)

= 6.01 x 10-6